# PCF Learned Sort: a Learning Augmented Sort Algorithm with $\mathcal{O}(n\log\log n)$ Expected Complexity

### v1

@@ 4.1 Results on Synthetic Datasets
We experimented with datasets created from the following four types of synthetic distributions: uniform distribution ($\mathsf{min}=0,\mathsf{max}=1$), normal distribution ($\mu=0,\sigma=1$), exponential distribution ($\lambda=1$), lognormal distribution ($\mu=0,\sigma=1$). We took $n=10^{3},\dots,10^{8}$ samples from each distribution independently, and meticulously counted the total number of basic operations to sort the sampled array. Here, the basic operations consists of four arithmetic operations, powers, comparisons, logical operations, assignments, and memory access. We chose this metric, which counts basic operations, to mitigate the environmental dependencies observed in other metrics, such as CPU instructions and CPU time, which are heavily influenced by compiler optimizations and the underlying hardware (this is the same idea as the metric selection in the experiment of  (Zeighami & Shahabi, 2023)). The parameters of PCF Learned Sort are set as in Theorem 3.5 and Theorem 3.6, $\alpha=\beta=\gamma=\delta=\lfloor n^{3/4}\rfloor$. As the “standard” sort algorithm used in PCF Learned Sort, we used quick sort, which is known to have an expected computational complexity of $\mathcal{O}(n\log n)$. We then compared the number of operations between PCF Learned Sort and the (plain) quick sort.
Figure 2 shows the number of operations divided by the length of input array, $n$. It shows the mean and standard deviation of the $10$ measurements for each condition. We see that our PCF Learned Sort has up to 2.8 times fewer operations than quick sort. Also, while the graph of quick sort is almost linear, the graph of PCF Learned Sort is almost flat. This suggests that PCF Learned Sort has a computational complexity much smaller than $\mathcal{O}(n\log n)$.
4.2 Results on Real Datasets
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 3: Number of operations to sort the real data array. Below each graph is a histogram that visualizes the distribution of each dataset.
We used four real datasets, NYC, Wiki, OSM, and Books, which are described below. NYC: pick-up datetimes in the yellow taxi trip records (Kristo, 2021). Wiki: Wikipedia article edit timestamps (Marcus et al., 2020). OSM: uniformly sampled OpenStreetMap locations represented as Google S2 CellIds (Marcus et al., 2020). Books: book sale popularity data from Amazon (Marcus et al., 2020). For each dataset, we randomly sample $n=10^{3},\dots,10^{8}$ elements, shuffle the sampled elements, and then use it as an input array to examine the performance of the algorithms. As in the case of the synthetic dataset, $\alpha=\beta=\gamma=\delta=\lfloor n^{3/4}\rfloor$ and quick sort is used as “standard” sort.
Figure 3 shows plots of the number of operations divided by the length of input array, $n$. Again, the mean and standard deviation of $10$ measurements are shown. We see that our PCF Learned Sort has up to 2.7 times fewer operations than quick sort (as in the case of the synthetic dataset), and that PCF Learned Sort has much less computational complexity than $\mathcal{O}(n\log n)$.
We can also see that the number of operations required for the OSM dataset is about $1.25$ times larger than that for the other datasets. Although not mentioned explicitly, the same tendency is observed in the experiments of existing Learned Sort studies (Kristo et al., 2020, 2021). Our theoretical analysis can explain this phenomenon from the perspective of $\sigma_{2}/\sigma_{1}$, in other words, the strength of the distributional bias. From Equation 11 and Equation 10, the larger $\sigma_{2}/\sigma_{1}$ is, the smaller $K$ is, and the larger (the upper bound of) the probability of bucketing failure. Therefore, the larger $\sigma_{2}/\sigma_{1}$ is, the more the computational complexity will move away from $\mathcal{O}(n\log\log n)$ and closer to the worst-case computational complexity, $\mathcal{O}(n\log n)$. In fact, the histograms of each dataset show that the OSM dataset has sharp peaks and many intervals with almost no data, indicating that it has a distribution with large $\sigma_{2}/\sigma_{1}$.
4.3 Confirmation of Lemma 3.4
@@ 4.2 Results on Real Datasets
Figure 4: Heatmap showing the experimental frequency of bucketing failure, i.e., $\exists j,|\bm{c}_{j}|>\delta$. The variables $a,b,c,d$, except those on the x- and y-axes, were set to $0.75$. The white dotted line represents the parameters that make the right side of Equation 10 equal to 0.5.
@@
Lemma 3.4 bounds the probability that a bucket of size greater than $\delta$ exists. This is an important lemma that allows us to guarantee the expected computational complexity of PCF Learned Sort. Here, we experimentally confirm that this upper bound is appropriate.
@@
We have experimented with $\alpha=\lfloor n^{a}\rfloor,\beta=\lfloor n^{b}\rfloor,\gamma=\lfloor n^{c}\rfloor,\delta=\lfloor n^{d}\rfloor$, varying $a,b,c,d$ from 0.05 to 0.95 at 0.05 intervals. For each $a,b,c,d$ setting, the following was repeated $100$ times: we took $n=10^{6}$ elements from the uniform distribution to form the input array and divided the array into $\gamma+1$ buckets by $\mathcal{M}_{\mathrm{PCF}}$, and checked whether or not $\exists j,|\bm{c}_{j}|>\delta$. Thus, for each $a,b,c,d\in\{0.05,0.10,\dots,0.95\}$, we obtained the experimental frequency at which bucketing “fails.”
@@
Heat maps in Figure 4 show the experimental frequency of bucketing failures when two of the $a,b,c,d$ parameters are fixed and the other two parameters are varied. The value of the two fixed variables are set to 0.75, e.g., in the upper left heap map of Figure 4 (horizontal axis is $a$ and vertical axis is $b$), $c=d=0.75$. The white dotted line represents the parameter so that the right side of Equation 10 is 0.5. That is, Lemma 3.4 asserts that “in the region upper right of the white dotted line, the probability of bucketing failure is less than 0.5.”
@@
We can see that the white dotted line is close to (or slightly to the upper right of) the actual bound of whether bucketing “succeeds” or “fails” more often. In other words, we can see that the theoretical upper bound from Lemma 3.4 agrees well (to some extent) with the actual probability. We can also confirm that, as Lemma 3.4 claims, the probability of bucketing failure is indeed small in the region upper right of the white line.
5 Limitation and Future Work
We proposed PCF Learned Sort, whose expected computational complexity can be proved to be $\mathcal{O}(n\log\log n)$ under the assumption that there exist $0<\sigma_{1}$ and $\sigma_{2}<\infty$ satisfying $\sigma_{1}<f(x)<\sigma_{2}$. Thus, our proof is not applicable to the case of a distribution where there exists $x\in\mathcal{D}$ such that $f(x)=0$ or $f(x)=\infty$. The development of a theory and/or algorithm that is applicable for a wider class of distributions is a future work.
Other future work may include incorporating a more advanced CDF approximation method with a theoretical guarantee into our Learned Sort algorithm. Using a different CDF model and a bucketing algorithm that satisfies the assumptions in Lemma 3.3, we may be able to develop an even more efficient sorting algorithm. In other words, a more accurate and faster CDF approximation method can lead to an improvement of the Learned Sort.
@@ 6 Conclusion
We proposed PCF Learned Sort and proved theoretically that its worst-case computational complexity is $\mathcal{O}(n\log n)$ (without assumptions) and its expected computational complexity is $\mathcal{O}(n\log\log n)$ under mild assumptions on the distribution. We then confirm this computational complexity experimentally on both synthetic and real data. Furthermore, we provide a theoretical and intuitive explanation for the slow speed of Learned Sort on some datasets. This is the first study to theoretically support the experimental success of Learned Sort and provides insight into why and when Learned Sort is fast.
Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Proofs
In this Appendix, we give the proofs omitted in the main paper. In Section A.1, we give the proof of Lemma 3.1, which is important for proving the worst-case complexity of PCF Learned Sort. Section A.2 and Section A.3 give proofs of Lemma 3.3 and Lemma 3.4, respectively, which are important for proving the expected computational complexity of PCF Learned Sort.
A.1 Proof of Lemma 3.1
@@
Let $P(n)$ be the worst-case complexity of our Learned Sort when using the model-based bucketing algorithm $\mathcal{M}$ as assumed in Lemma 3.1 and $\delta=\lfloor n^{d}\rfloor$. Let $S(n)$ be the worst-case complexity of the “standard” sort algorithm and $R(n)$ be the worst-case complexity of model-based bucketing (including model training and inferences). Since $S(n)=\mathcal{O}(n\log n)$ and the “standard” sorting algorithm terminates after a finite number of operations, $R(n)=\mathcal{O}(n)$, and $\gamma+1=\mathcal{O}(n)$,
@@
First, for $n<\max(n_{2},n_{3},\tau)\eqcolon n_{0}$, there exists a constant $C~{}(>0)$ such that $P(n)\leq C$. That is, for $n<n_{0}$, our Learned Sort terminates in a finite number of operations. This is because, since $\delta<n$, the bucket will either be smaller than the original array length $n$, or the bucket will be immediately sorted by the “standard” sort algorithm.
Next, assume that there exists a constant $C~{}(>0)$ and $l~{}(>0)$ such that $P(n)\leq C+ln\log n$ for all $n<k$, where $k$ is an integer such that $k\geq n_{0}$. Let $\mathcal{S}_{\gamma,k}$ be the set consisting of all $(\gamma+1)$-dimensional vectors of positive integers whose sum is $k$, i.e., $\mathcal{S}_{\gamma,k}\coloneq\left\{\bm{s}\in\mathbb{Z}_{\geq 0}^{\gamma}\mid\sum_{i=1}^{\gamma+1}s_{i}=k\right\}$. Then,
$\begin{split}P(k)&\leq R(k)+\max_{\bm{s}\in\mathcal{S}}\sum_{i=1}^{\gamma+1}\left\{\mathbbm{1}[s_{i}\geq\lfloor k^{d}\rfloor]\cdot S(s_{i})+\mathbbm{1}[s_{i}<\lfloor k^{d}\rfloor]\cdot P(s_{i})\right\}\\ &\leq l_{2}k+\max_{\bm{s}\in\mathcal{S}}\sum_{i=1}^{\gamma+1}\left\{\mathbbm{1}[s_{i}\geq\lfloor k^{d}\rfloor]\cdot\left(C_{1}+l_{1}s_{i}\log s_{i}\right)+\mathbbm{1}[s_{i}<\lfloor k^{d}\rfloor]\cdot\left(C+ls_{i}\log s_{i}\right)\right\}\\ &\leq l_{2}k+\max_{\bm{s}\in\mathcal{S}}\sum_{i=1}^{\gamma+1}\left\{\left(C_{1}+l_{1}s_{i}\log s_{i}\right)+\left(C+ls_{i}\log k^{d}\right)\right\}\\ &\leq l_{2}k+C_{1}(\gamma+1)+l_{1}k\log k+C(\gamma+1)+lk\log k^{d}\\ &\leq l_{2}k+C_{1}l_{3}k+l_{1}k\log k+Cl_{3}k+lkd\log k\\ &\leq\left\{l_{2}+C_{1}l_{3}+l_{1}+Cl_{3}-l(1-d)\right\}k\log k+(C+lk\log k).\end{split}$ (16)
@@
$\frac{l_{2}+C_{1}l_{3}+l_{1}+Cl_{3}}{1-d}\leq l,$ (17)
then $P(k)\leq C+lk\log k$ (note that the left side of the following equation is a constant independent of $k$).
Hence, by mathematical induction, it is proved that there exists a constant $C~{}(>0)$ and $l~{}(>0)$ such that $P(n)\leq C+ln\log n$ for all $n\in\mathbb{N}$. ∎
A.2 Proof of Lemma 3.3
@@
The proof approach is the same as in Lemma 3.1, but in Lemma 3.3 the “expected” computational complexity is bounded. The following two randomnesses are considered to compute the “expected” computational complexity: (i) the randomness with which $n$ elements are independently sampled according to the probability density function $f(x)$ in the process of forming the input array $\bm{x}$, and (ii) the randomness of the PCF Learned Sort algorithm sampling $\alpha$ elements from the input array $\bm{x}$ for training the PCF.
Let $T(n)$ be the expected complexity of our Learned Sort when using the model-based bucketing algorithm $\mathcal{M}$ as assumed in Lemma 3.3 and $\delta=\lfloor n^{d}\rfloor$. Let $S(n)$ be the expected complexity of the “standard” sort algorithm and $R(n)$ be the expected complexity of model-based bucketing (including model training and inferences). Since $S(n)=\mathcal{O}(n\log n)$ and the “standard” sorting algorithm terminates after a finite number of operations, $R(n)=\mathcal{O}(n)$, $\Pr[\exists j,|\bm{c}_{j}|\geq\lfloor n^{d}\rfloor]=\mathcal{O}(1/\log n)$, and $\gamma+1=\mathcal{O}(n)$,
$\exists~{}C_{1},l_{1}~{}(>0),$ $~{}~{}n\geq 0\Rightarrow S(n)\leq C_{1}+l_{1}n\log n,$ (18)
@@
First, for $n<\max(n_{2},n_{3},n_{4},\tau)\eqcolon n_{0}$, there exists a constant $C~{}(>0)$ such that $T(n)\leq C$. That is, for $n<n_{0}$, our Learned Sort terminates in a finite number of operations.
Next, assume that there exists a constant $C~{}(>0)$ and $l~{}(>0)$ such that $T(n)\leq C+ln\log\log n$ for all $n<k$, where $k$ is an integer such that $k\geq n_{0}$. Then, from $k\geq n_{2}$,
@@
then $T(k)\leq C+lk\log\log k$ (note that the left side of Equation 26 is a constant independent of $k$).
Hence, by mathematical induction, it is proved that there exists a constant $C~{}(>0)$ and $l~{}(>0)$ such that $T(n)\leq C+ln\log\log n$ for all $n\in\mathbb{N}$. ∎
A.3 Proof of Lemma 3.4
To prove Lemma 3.4, we first present the following lemma.
@@
Also, define the set $\mathcal{S}_{r},\mathcal{T}_{r}$ as follows ($r=1,\dots,n$):
$\mathcal{S}_{r}=\{k\mid e_{\max(1,r-\delta/2)}+\Delta<e_{k}\leq e_{r}-\Delta\},~{}~{}~{}~{}~{}~{}~{}~{}\mathcal{T}_{r}=\{k\mid e_{r}+\Delta\leq e_{k}<e_{\min(r+\delta/2,n)}-\Delta\}.$ (27)

### v3

@@ 4.1 Computational Complexity of PCF Learned Sort
We meticulously counted the total number of basic operations for sorting the input array to observe the computational complexity of each sorting algorithm. Here, the basic operations consist of four arithmetic operations, powers, comparisons, logical operations, assignments, and memory access. We chose this metric, which counts basic operations, to mitigate the environmental dependencies observed in other metrics, such as CPU instructions and CPU time, which are heavily influenced by compiler optimizations and the underlying hardware. This is the same idea as the metric selection in the experiment of  (zeighami2023distribution).
The parameters of PCF Learned Sort are set as in Theorems˜3.5 and 3.6, $\alpha=\beta=\gamma=\delta=\lfloor n^{3/4}\rfloor$ and $\tau=100$. As the standard sort, we used QuickSort, which has an expected complexity of $\mathcal{O}(n\log n)$. For ease of implementation and straightforward measurement of comparison counts, we chose QuickSort. We compared our PCF Learned Sort against (plain) QuickSort, radix sort, and Learned Sort 2.0 (kristo2021defeating-github). Learned Sort 2.0 was selected as the learned-method baseline because it allows relatively simple counting of operations (a broader comparison appears in the sorting time experiments of Section˜4.3).
Figure 2: Number of operations to sort the array. Below each graph is a histogram visualizing the distribution of each dataset. The standard deviation of the $10$ measurements is represented by the shaded area. Our PCF Learned Sort consistently achieves a complexity lower than $\mathcal{O}(n\log n)$, while Learned Sort 2.0 (kristo2021defeating), which has $\mathcal{O}(n^{2})$ worst-case complexity, occasionally requires huge operations.
Figure˜2 shows the number of operations divided by the length of the input array, $n$. Each point represents the average over 10 runs, with the shaded region indicating the standard deviation. Note that the horizontal axis is logarithmic. As a result, the curve for QuickSort (with $\mathcal{O}(n\log n)$ complexity) appears approximately linear in this plot for synthetic datasets and real-world datasets with few duplicates (i.e., NYC [Pickup], Wiki, OSM, Books, and Face). In contrast, the curve for PCF Learned Sort is nearly flat, suggesting a complexity significantly lower than $\mathcal{O}(n\log n)$. PCF Learned Sort consistently requires the fewest operations across all conditions, achieving up to 2.8 times fewer operations than QuickSort. These results confirm not only our theoretical analysis but also the robustness of PCF Learned Sort under assumption-violating scenarios.
For datasets with many duplicates, QuickSort, Learned Sort 2.0 (kristo2021defeating), and PCF Learned Sort all exhibit relatively few operations. This behavior arises because highly duplicated datasets often result in buckets containing only one distinct value, allowing the algorithms to terminate early. Even under such conditions, PCF Learned Sort performs fewer operations than the other algorithms.
The curve for radix sort is also flat but lies consistently above that of PCF Learned Sort. This is due to the difference in partitioning strategies. While the radix sort partitions at a fixed granularity, our PCF Learned Sort performs partitioning using intervals that are adaptively set by the learning model.
In particular, on the SOF [Temperature] dataset with medium input sizes ($2\times 10^{5}<n\leq 2\times 10^{7}$), Learned Sort 2.0 incurs extremely high costs: while all other methods keep the number of operations per $n$ below 100, it can exceed 50,00050{,}000. This shows that Learned Sort methods without worst-case guarantees may suffer from pathological overhead. By contrast, our method maintains both expected and worst-case guarantees, ensuring robust performance across diverse datasets.
4.2 Confirmation of Lemma˜3.4
@@
Figure 3: Heatmap showing the empirical frequency of bucketing failure, i.e., $\exists j,|\bm{c}_{j}|>\delta$. The variables $a,b,c,d$, except those on the x- and y-axes, were set to $0.75$. The white dotted line represents the parameters that make the right side of Equation˜1 equal to 0.5. The close alignment between this white dotted line and the actual success/fail boundery suggests the theoretical bound by is Equation˜1 reasonably tight.
@@
Lemma˜3.4 bounds the probability that a bucket of size greater than $\delta$ exists. This is an important lemma that allows us to guarantee the expected computational complexity of PCF Learned Sort. Here, we empirically confirm that this upper bound is appropriate.
@@
We have experimented with $\alpha=\lfloor n^{a}\rfloor,\beta=\lfloor n^{b}\rfloor,\gamma=\lfloor n^{c}\rfloor,\delta=\lfloor n^{d}\rfloor$, varying $a,b,c,d$ from 0.05 to 0.95 at 0.05 intervals. For each $a,b,c,d$ setting, the following was repeated $100$ times: we took $n=10^{6}$ elements from the uniform distribution to form the input array and divided the array into $\gamma+1$ buckets by $\mathcal{M}_{\mathrm{PCF}}$, and checked whether or not $\exists j,|\bm{c}_{j}|>\delta$. Thus, for each $a,b,c,d\in\{0.05,0.10,\dots,0.95\}$, we obtained the empirical frequency at which bucketing “fails.”
@@
Heat maps in Figure˜3 show the empirical frequency of bucketing failures when two of the $a,b,c,d$ parameters are fixed, and the other two parameters are varied. The values of the two fixed variables are set to 0.75, e.g., in the upper left heap map of Figure˜3 (horizontal axis is $a$ and vertical axis is $b$), $c=d=0.75$. The white dotted line represents the parameter so that the right side of Equation˜1 is 0.5. That is, Lemma˜3.4 asserts that “in the region upper right of the white dotted line, the probability of bucketing failure is less than 0.5.”
@@
We observe that the white dotted line is close to (or slightly to the upper right of) the actual bound of whether bucketing “succeeds” or “fails” more often. This suggests that the theoretical upper bound from Lemma˜3.4 agrees well (to some extent) with the actual probability. We can also confirm that, as Lemma˜3.4 claims, the probability of bucketing failure is indeed small in the region upper right of the white line.
4.3 Experiments on Sorting Time
We empirically compare the sorting time of our PCF Learned Sort against several baselines. Note that the metric used in Section˜4.1, the number of operations, does not change depending on the machine or compilation method, but the sorting time does. As the standard sort used in PCF Learned Sort, we used std::sort, which has a worst-case complexity of $\mathcal{O}(n\log n)$. We compared our PCF Learned Sort with std::sort, radix sort, boost::sort::spreadsort::float_sort (Boost C++ implementation of Spreadsort (ross2002spreadsort)), and Learned Sort 2.0 (kristo2021defeating-github). In addition, we evaluated more recent state-of-the-art learned sorting algorithms (ferragina2025balancedsort)—Balanced Learned Sort (BLS), Unbalanced Learned Sort (ULS), and Learned Sort 2.1—as well as IS4o (axtmann2017), one of the state-of-the-art non-learned sequential sample sort algorithms.
Figure 4: Time to sort the array. The standard deviation of the $10$ measurements is represented by the shaded area. Our PCF Learned Sort is significantly faster than std::sort, and more importantly, it maintains robust performance across all datasets, whereas other learned sorts without worst-case guarantees can suffer catastrophic slowdowns, as seen with Learned Sort 2.0 on the SOF [Temperature] dataset.
Figure˜4 shows the sorting time divided by $n$. It shows the mean and standard deviation of the $10$ measurements for each condition. Note that the horizontal axis is logarithmic, and therefore the curve of std::sort, which has a complexity of $\mathcal{O}(n\log n)$, is almost linear in this plot for synthetic datasets and real-world datasets with few duplicates. In contrast, our PCF Learned Sort graph shows a relatively slow increase, suggesting that PCF Learned Sort has a complexity much smaller than $\mathcal{O}(n\log n)$. We see that our PCF Learned Sort is up to 2.5 times faster than std::sort. Furthermore, we find that for relatively large $n$ ($10^{6}$), our PCF Learned Sort usually outperforms not only radix sort but also Spreadsort, an algorithm that cleverly incorporates the advantage of comparison sort into radix sort.
The figure also shows that highly optimized methods like IS4o and other learned sorts often outperform PCF Learned Sort in average speed. This is because these methods are highly optimized implementations that consider factors like cache efficiency, whereas our implementation prioritizes providing rigorous theoretical guarantees. However, this speed comes at the cost of robustness: learned sorts without worst-case guarantees can suffer catastrophic slowdowns. For example, on the SOF [Temperature] dataset with $n=2\times 10^{7}$, Learned Sort 2.0 took up to 326.4 seconds, while our method consistently finished in 0.45 seconds. Moreover, as shown in our adversarial analysis (Appendix˜C), other recent learned sorts (BLS, ULS, Learned Sort 2.1) also exhibit vulnerabilities. These results highlight the practical value of our theoretical guarantees.
To better understand runtime behavior, we profiled PCF Learned Sort across its stages (Appendix˜D). The results show that while the training stage is consistently lightweight across datasets and input sizes, the bucketing stage often becomes the bottleneck. The dominant cost varies depending on both $n$ and the dataset characteristics, suggesting opportunities for further optimization.
5 Discussion
Comparison with (zeighami2024theoretical).
Most recently, a concurrent work (zeighami2024theoretical) introduced a theoretical framework for learned database operations, including sorting, indexing, and cardinality estimation. A key strength of their approach is the formal definition of “distribution learnability” and its applicability to a broad class of distributions, including those subject to distribution shifts. Using this framework, they developed a Learned Sort algorithm with an expected running time of $O(n\log\log n)$ under certain distributional assumptions. While their approach employs a bucketing-based sorting algorithm similar to ours, there are several key differences between their method and ours.
First, their algorithm relies on detailed prior knowledge about the distribution; it explicitly requires a parameter $\varkappa_{2}$, which represents the “learning possibility” of the distribution (see Definition 3.2 in (zeighami2024theoretical) for details). Since $\varkappa_{2}$ depends on $\rho_{1}$ and $\rho_{2}$, their algorithm cannot be executed without knowing the values of $\rho_{1}$ and $\rho_{2}$ (or at least their lower and upper bounds). In contrast, our algorithm does not require these values, making it more widely applicable.
Second, they do not provide experimental results. Since estimating $\rho_{1}$ and $\rho_{2}$ from real data is challenging, empirical evaluation of their method is difficult. On the other hand, since our algorithm does not depend on these specific values, it is easy to implement and evaluate experimentally.
Finally, their algorithm lacks a worst-case complexity guarantee. In particular, there are cases where the algorithm may not terminate, making it impossible to give an upper bound on its worst-case complexity. In contrast, we provide a formal worst-case complexity guarantee.
Limitations of the Current Theoretical Framework.
Our proof of the $\mathcal{O}(n\log\log n)$ expected complexity (Theorem˜3.6), does not extend to distributions with $f(x)=0$ or $\infty$. A similar limitation is observed in Learned Indexes (zeighami2023distribution), and extending the theory to cover a broader class of distributions remains an open direction in both Learned Indexes and Learned Sorts contexts. One promising direction is to integrate more advanced CDF approximation methods with theoretical guarantees into our Learned Sort algorithm. By adopting a refined CDF model and a bucketing algorithm that satisfies the conditions of Lemma˜3.3, it may be possible to achieve stronger theoretical guarantees.
Implementation Considerations and Optimizations.
A straightforward implementation of PCF Learned Sort is not in-place, as it requires an auxiliary buffer for buckets nearly as large as the input. Engineering techniques from highly optimized sample-sort implementations (axtmann2017; axtmann2022engineering) suggest in-place variants without changing the asymptotic structure of our algorithm. Another important aspect is real-world efficiency: optimizing memory access patterns to enhance cache efficiency, reducing cache misses, and leveraging cache-aware data layouts could improve empirical performance without compromising theoretical guarantees. Dynamically tuning parameters such as $\alpha$, $\beta$, $\gamma$, and $\delta$ based on the input distribution may yield further performance gains while maintaining guarantees.
Parallelization Potential.
Although our analysis targets the sequential setting, the structure of PCF Learned Sort admits parallelization at several stages: training the CDF model, computing bucket IDs, scattering into buckets, and sorting buckets are all amenable to data parallelism. This observation is consistent with prior engineering work that integrates learned partitioning with high-performance (in-place) sample-sort pipelines and parallel learned-sorting frameworks (axtmann2017; carvalho2022towards; carvalho2023learnedsort). These results complement our theoretical analysis and suggest opportunities for developing parallel learned-sorting algorithms with provable guarantees.
@@ 6 Conclusion
We proposed PCF Learned Sort, which provides guarantees on both its expected and worst-case complexities. We then confirmed these computational complexities empirically on both synthetic and real data. This is the first study to support the empirical success of Learned Sort theoretically and provides insight into why Learned Sort is fast.
@@ Appendix A Proofs
Here, we give the proofs omitted in the main paper. In Section˜A.1, we give the proof of Lemma˜3.1, which is important for proving the worst-case complexity of PCF Learned Sort. Section˜A.2 and Section˜A.3 give proofs of Lemma˜3.3 and Lemma˜3.4, respectively, which are important for proving the expected computational complexity of PCF Learned Sort. The quantization-aware version of Lemma˜3.4 and Theorem˜3.6 is defined and the proof is given in the Section˜A.4. Finally, we provide the proofs of the theorems guaranteeing the complexity of Spline Learned Sort in Section˜A.5.
A.1 Proof of Lemma˜3.1
@@
Let $P(n)$ be the worst-case complexity of our Learned Sort when using the model-based bucketing algorithm $\mathcal{M}$ as assumed in Lemma˜3.1 and $\delta=\lfloor n^{d}\rfloor$. Let $S(n)$ be the worst-case complexity of the standard sort and $R(n)$ be the worst-case complexity of model-based bucketing (including model training and inferences). Since $S(n)=\mathcal{O}(nU(n))$ and the standard sort terminates after a finite number of operations,
$\exists~C_{1},l_{1}~(>0),~~n\geq 0\;\;\Rightarrow\;\;S(n)\leq C_{1}+l_{1}nU(n).$ (3)
Since $R(n)=\mathcal{O}(n)$ and $\gamma+1=\mathcal{O}(n)$,
@@
In the following, we prove $P(n)=\mathcal{O}(nU(n)+n\log\log n)$ by mathematical induction.
First, for $n<\max(n_{2},n_{3},\tau)\eqcolon n_{0}$, there exists a constant C(>0)C~(>0) such that $P(n)\leq C$. That is, for $n<n_{0}$, our Learned Sort terminates in a finite number of operations. This is because, since $\delta<n$, the bucket will either be smaller than the original array length $n$, or the bucket will be immediately sorted by the standard sort.
Next, assume that there exists a constant C(>0)C~(>0) and l(>0)l~(>0) such that $P(n)\leq C+l(nU(n)+n\log\log n)$ for all $n<k$, where $k$ is an integer such that $k\geq n_{0}$. Without loss of generality, we assume that $C\geq C_{1}$ and $l\geq l_{1}$. Let $\mathcal{S}_{\gamma,k}$ be the set consisting of all $(\gamma+1)$-dimensional vectors of positive integers whose sum is $k$, i.e., $\mathcal{S}_{\gamma,k}\coloneq\left\{\bm{s}\in\mathbb{Z}_{\geq 0}^{\gamma}\mid\sum_{i=1}^{\gamma+1}s_{i}=k\right\}$. Then,
$\begin{split}P(k)&\leq R(k)+\max_{\bm{s}\in\mathcal{S}_{\gamma,k}}\sum_{i=1}^{\gamma+1}\left\{\mathbbm{1}[s_{i}\geq\lfloor k^{d}\rfloor]\cdot S(s_{i})+\mathbbm{1}[s_{i}<\lfloor k^{d}\rfloor]\cdot P(s_{i})\right\}\\ &\leq l_{2}k+\max_{\bm{s}\in\mathcal{S}_{\gamma,k}}\sum_{i=1}^{\gamma+1}\left\{\mathbbm{1}[s_{i}\geq\lfloor k^{d}\rfloor]\cdot\left(C_{1}+l_{1}s_{i}U(s_{i})\right)+\mathbbm{1}[s_{i}<\lfloor k^{d}\rfloor]\cdot\left(C+l(s_{i}U(s_{i})+s_{i}\log\log s_{i})\right)\right\}\\ &\leq l_{2}k+\max_{\bm{s}\in\mathcal{S}_{\gamma,k}}\sum_{i=1}^{\gamma+1}\left(C+ls_{i}U(s_{i})+ls_{i}\log\log k^{d}\right)\\ &\leq l_{2}k+C(\gamma+1)+lkU(k)+lk\log\log k^{d}\\ &\leq l_{2}k+Cl_{3}k+lkU(k)+lk\log\log k+lk\log d\\ &\leq\left(l_{2}+Cl_{3}-l\log\frac{1}{d}\right)k+(C+l(kU(k)+k\log\log k)).\end{split}$ (6)
@@
$\frac{l_{2}+Cl_{3}}{\log\frac{1}{d}}\leq l,$ (7)
then $P(k)\leq C+l(kU(k)+k\log\log k)$ (note that the left side of Equation˜7 is a constant independent of $k$).
Hence, by mathematical induction, it is proved that there exists a constant C(>0)C~(>0) and l(>0)l~(>0) such that $P(n)\leq C+l(kU(k)+lk\log\log k)$ for all $n\in\mathbb{N}$. ∎
A.2 Proof of Lemma˜3.3
@@
The proof approach is the same as in Lemma˜3.1, but in Lemma˜3.3, the “expected” computational complexity is bounded. The following two randomnesses are considered for computing the “expected” computational complexity: (i) the randomness with which $n$ elements are independently sampled according to the probability density function $f(x)$ in the process of forming the input array $\bm{x}$, and (ii) the randomness of the PCF Learned Sort algorithm sampling $\alpha$ elements from the input array $\bm{x}$ for training the PCF.
Let $T(n)$ be the expected complexity of our Learned Sort when using the model-based bucketing algorithm $\mathcal{M}$ as assumed in Lemma˜3.3 and $\delta=\lfloor n^{d}\rfloor$. Let $S(n)$ be the expected complexity of the standard sort and $R(n)$ be the expected complexity of model-based bucketing (including model training and inferences). Since $S(n)=\mathcal{O}(n\log n)$ and the standard sort terminates after a finite number of operations,
Since $R(n)=\mathcal{O}(n)$, $\Pr[\exists j,|\bm{c}_{j}|\geq\lfloor n^{d}\rfloor]=\mathcal{O}(1/\log n)$, and $\gamma+1=\mathcal{O}(n)$,
@@
First, for $n<\max(n_{2},n_{3},n_{4},\tau)\eqcolon n_{0}$, there exists a constant C(>0)C~(>0) such that $T(n)\leq C$. That is, for $n<n_{0}$, our Learned Sort terminates in a finite number of operations.
Next, assume that there exists a constant C(>0)C~(>0) and l(>0)l~(>0) such that $T(n)\leq C+ln\log\log n$ for all $n<k$, where $k$ is an integer such that $k\geq n_{0}$. Then, from $k\geq n_{2}$,
@@
then $T(k)\leq C+lk\log\log k$ (note that the left side of Equation˜16 is a constant independent of $k$).
Hence, by mathematical induction, it is proved that there exists a constant C(>0)C~(>0) and l(>0)l~(>0) such that $T(n)\leq C+ln\log\log n$ for all $n\in\mathbb{N}$. ∎
A.3 Proof of Lemma˜3.4
We first present the following lemma to prove Lemma˜3.4.
@@
Also, define the set $\mathcal{S}_{r}$ and $\mathcal{T}_{r}$ as follows ($r=1,\dots,n$):
$\mathcal{S}_{r}=\{k\mid e_{\max(1,r-\delta/2)}+\Delta<e_{k}\leq e_{r}-\Delta\},~~~~~\mathcal{T}_{r}=\{k\mid e_{r}+\Delta\leq e_{k}<e_{\min(r+\delta/2,n)}-\Delta\}.$ (17)